### v1

# Assessing Vulnerability in Smart Contracts: The Role of Code Complexity Metrics in Security Analysis

Abstract
Codes with specific characteristics are more exposed to security vulnerabilities. Studies have revealed that codes that do not adhere to best practices are more challenging to verify and maintain, increasing the likelihood of unnoticed or unintentionally introduced vulnerabilities. Given the crucial role of smart contracts in blockchain systems, ensuring their security and conducting thorough vulnerability analysis is critical. This study investigates the use of code complexity metrics as indicators of vulnerable code in Solidity smart contracts. We highlight the significance of complexity metrics as valuable complementary features for vulnerability assessment and provide insights into the individual power of each metric. By analyzing 21 complexity metrics, we explored their interrelation, association with vulnerability, discriminative power, and mean values in vulnerable versus neutral codes. The results revealed some high correlations and potential redundancies among certain metrics, but weak correlations between each independent metric and vulnerability. Nevertheless, we found that all metrics can effectively discriminate between vulnerable and neutral codes, and most complexity metrics, except for three, exhibited higher values in vulnerable codes.
Index Terms:
Smart contract, Software Metrics, Complexity Metrics, Vulnerability detection, Statistical Analysis
I Introduction
Ethereum, referenced as [1], is a decentralized computing platform and operating system based on blockchain technology. It is an open-source platform that caters to diverse business scenarios by offering Solidity, a Turing-complete programming language, and the Ethereum Virtual Machine (EVM). Developers can utilize these tools to deploy smart contracts [2] on the Ethereum blockchain. smart contracts find widespread applications in various industries, including decentralized finance (DeFi) within financial services [3], infrastructure development [4], the internet of things (IoT) [5], the gaming industry [3], healthcare [6, 7], the metaverse [8], and numerous other sectors [3]. In contrast to traditional applications, smart contracts are executed automatically in shortcodes, which safeguards them against manipulation during the execution process. Open-source availability is a common characteristic of widely-used smart contracts, infusing trust among users regarding the integrity of the code. Since smart contracts operate on distributed nodes within the blockchain, once a contract is executed and deployed in a block, it becomes quite challenging to upgrade or modify the contract [9]. This is due to the inherent immutability of the blockchain, where tampering with any block is virtually impossible. Furthermore, smart contracts have the capability to hold funds within their associated blocks on the blockchain. These unique properties of smart contracts make them an attractive target for hackers. While the underlying blockchain technology itself is highly secure and resistant to hacking attempts, smart contracts can still be vulnerable to attacks. This vulnerability often stems from factors such as programmer negligence or errors in the code. Therefore, implementing robust vulnerability detection mechanisms becomes essential to identify and address any potential security flaws in smart contracts prior to their deployment [10, 11]. The security of smart contracts is of utmost importance, as evidenced by the significant amount of money lost in real-world smart contract vulnerabilities each month. A notable example is the attack on the Decentralized Autonomous Organization (DAO) in 2016, where 3.6 million Ether was stolen using a simple recursive vulnerability [12]. Additionally, according to blockchaingroup.io  [^1] https://blockchaingroup.io/compliance-and-regulation/top-10-crypto-losses-of-2024-hacks-frauds-and-exploits/, the top 10 crypto hacks in 2023 alone resulted in the theft of 1.06 billion US dollars by hackers. Table I provides an overview of the amounts stolen in each smart contract hack. These staggering numbers serve as evidence of the critical importance of ensuring the security of smart contracts.
The rest of the paper is structured as follows. Section two presents our contribution and outlines the research questions we aim to answer. Section three provides background and related work, offering the necessary knowledge to understand this study and reviewing previous work on the topic. Section four details each complexity metric used in this study. Section five describes the dataset utilized. Section six, the experimental setup, explains the techniques and analysis methods employed to answer the research questions. Section seven presents the results of our analysis, addressing each research question in detail and analyzing the answers. Section eight, the discussion, explores the limitations of this study and suggests future directions. Lastly, section nine concludes the paper, summarizing the answers to the research questions and the overall findings.
TABLE I: Top 10 smart contract hacks and losses in 2024, according to blockchaingroup.io
smart contract Loss(\$)
UwU Lend 19.3M
Indodax 22M
BinX Exchange 26M
Penpie 27M
Hedgey Finance 44M
BtcTurk Hot Wallet 55M
DMM Bitcoin Private Key 125M
Orbit Chain Bridge 197M
WazirX 230M
PlayDapp 290M
Total 1.03B
II Contribution & Research Questions
The primary focus of this study is to investigate the impact of complexity metrics on vulnerabilities present in Solidity smart contracts. While existing static analysis tools are designed to detect specific vulnerability patterns, complexity metrics offer a complementary approach to identifying vulnerable areas within smart contracts. Our specific research objectives are as follows:
  • Determine whether there are noticeable differences in complexity metrics between vulnerable and non-vulnerable functions. If such differences exist, we aim to identify which complexity metrics are most effective in distinguishing between the two.
  • Explore the feasibility of using complexity metrics to predict vulnerable contracts among all contracts within a smart contract ecosystem.
This investigation aims to assess whether complexity metrics can serve as indicators for identifying vulnerable code locations. Furthermore, we aim to highlight the significance of complexity metrics as valuable complementary features for vulnerability assessment and provide insights into the individual power of each metric. Thus, we address four research questions related to software metrics and vulnerability detection:
  • RQ1: Are there correlations among complexity metrics? Our objective is to identify potential redundancies or inconsistencies among the various complexity metrics available.
  • RQ2: Is there a correlation between each complexity metric and the existence of vulnerability? We aim to determine whether certain complexity metrics are associated with the presence of vulnerabilities in individual contracts.
  • RQ3: Are the metrics different in vulnerable and neutral codes? We seek to investigate whether complexity metrics can effectively distinguish between functions with reported vulnerabilities and those without reported vulnerabilities.
  • RQ4: How different are the metrics in vulnerable codes compared to neutral ones? We intend to examine whether higher complexity in code is linked to a higher likelihood of vulnerability occurrence. This approach enables a more targeted and efficient allocation of resources toward securing smart contracts.
The graphical representation of this paper’s contributions and a brief outline of how the research questions are addressed can be found in Figure 1.
Figure 1: Contribution and outline of this paper
III Background & Related Work
Researchers have been actively proposing various methods and tools to identify security vulnerabilities in smart contracts. These approaches include static and dynamic code analysis, as well as formal methods [13] and Machine learning.
III-A Static Analysis
Static code analysis involves examining the source code of a smart contract without executing it, aiming to detect potential vulnerabilities through code inspection. Slither [14] is a Python 3-based open-source static analysis framework that relies on the Solidity compiler (Solc) as its dependency. It offers swift, reliable, precise, and comprehensive information about smart contracts. Slither makes use of SlithIR, an intermediate representation designed for practical analysis of Solidity code. In addition to bug detection, it provides suggestions for code optimizations and enhances the comprehension of smart contract code [15]. In addition to Slither, there are several other noteworthy static analysis tools for smart contracts, such as SolAnalyser [16] and EtherTrust [17].
III-B Dynamic Analysis
Dynamic code analysis, on the other hand, involves analyzing the behavior of a smart contract during execution to identify any security flaws.
III-C Machine Learning
Machine learning techniques have proven to be valuable in the detection of software vulnerabilities [18]. By analyzing vast amounts of data and identifying patterns or anomalies, machine learning algorithms can effectively pinpoint potential weaknesses in software. This approach serves as a complementary method, augmenting other analysis approaches by providing automated analysis capabilities and enabling scalability in the vulnerability detection process. Machine learning’s ability to learn from data and identify subtle patterns makes it a promising tool for enhancing the efficiency and effectiveness of software vulnerability detection.
A recent machine learning (ML) model, referred to as MODNN [19], demonstrated success in utilizing a publicly available dataset to evaluate its approach. By leveraging this dataset, the MODNN model showcased its efficacy in analyzing and addressing vulnerable smart contracts. The proposal introduces the Multiple-Objective Detection Neural Network (MODNN) as a scalable smart contract vulnerability detection tool, leveraging the ML MOD (Multiple-Objective Detection) method. The aim is to enhance the security of cryptocurrency systems. A novel approach is also presented, focusing on extracting implicit features from operation sequences to support the scalability of the MODNN model. By combining these advancements, the research aims to provide a more efficient and effective solution for detecting vulnerabilities in smart contracts within cryptocurrency systems.
The researchers in DR-GCN [20] investigate the use of graph neural networks (GNN) and expert knowledge for detecting vulnerabilities in smart contracts. They focus on capturing the extensive control-flow and data-flow semantics of the source code by representing it as a contract graph. To enhance the interpretability of the graph, they introduce a node elimination phase to normalize it and identify critical nodes. Subsequently, they propose a unique temporal message propagation network that extracts graph features from the normalized graph. These graph features are then combined with expert patterns that have been specifically designed for vulnerability detection. The approach outlined in the paper aims to leverage both GNN and expert knowledge to improve the accuracy and effectiveness of vulnerability detection in smart contracts.
In Soliaudit, a study conducted by Liao et al. [21], machine learning and fuzz testing techniques are employed for assessing vulnerabilities in smart contracts. SoliAudit utilizes machine learning technology, specifically using Solidity machine code as learning features, to verify 13 different types of vulnerabilities. The researchers also develop a gray-box fuzz testing mechanism comprising a fuzzer contract and a simulated blockchain environment for online transaction verification. Notably, SoliAudit distinguishes itself from previous research systems by its ability to detect vulnerabilities without relying on expert knowledge or predefined patterns.
Eth2Vec [22] is a machine-learning-based static analysis tool designed to detect vulnerabilities in smart contracts. It operates by automatically learning the features of vulnerable Ethereum Virtual Machine (EVM) bytecodes using a neural network for natural language processing. This approach enables Eth2Vec to acquire tacit knowledge and expertise in identifying vulnerabilities. By leveraging the learned knowledge, Eth2Vec can effectively detect vulnerabilities in smart contracts by comparing the code similarity between the target EVM bytecodes and the bytecodes it has already learned. These successful evaluations highlight the potential of ML models to harness public datasets, enabling researchers to refine and enhance their approaches to tackle complex problems. These combined efforts contribute to the ongoing development of effective approaches for identifying and mitigating security vulnerabilities in smart contracts.
While machine learning models have made significant steps, they still face challenges in effectively detecting vulnerabilities within smart contracts. To maintain their capabilities, the incorporation of additional features is crucial. Notably, the potential of software metrics, particularly complexity metrics, has yet to be thoroughly explored in terms of their effectiveness in identifying vulnerabilities in the domain of smart contracts.
III-D Complexity and Vulnerability
Defining program complexity is a crucial step in understanding software systems. According to Basili, complexity can be seen as a measure of the resources utilized by a system while interacting with a piece of software to accomplish a specific task [23]. The interpretation of complexity depends on the interacting system, whether it is a computer or a programmer. In the context of a computer, complexity is determined by the execution time and storage requirements for computation. On the other hand, when considering the interaction between a program and a programmer, complexity is characterized by the challenges involved in tasks such as coding, debugging, testing, and modifying the software. The term “software complexity” often refers to the interaction between a program and a programmer engaged in programming activities [24].
Security experts assert that complexity poses a significant threat to security [25, 26]. Complex code can introduce subtle vulnerabilities that are challenging to detect and diagnose [26], thereby offering attackers greater opportunities for exploitation. Furthermore, complex code presents difficulties in comprehension, maintenance, and testing. Consequently, complex code is more prone to faults compared to simpler code. As attackers exploit these faults in a program, complex code becomes more vulnerable than its simpler counterparts. Until now, the use of complexity metrics as indicators of syntactic and semantic information extracted from smart contracts has not been extensively explored [27]. However, the effectiveness of complexity metrics in identifying vulnerable code has been extensively studied and validated in various applications and programming languages [28]. Researchers have leveraged techniques like machine learning and statistical models to demonstrate the correlation between code complexity and the presence of vulnerabilities.
In a study conducted by Shin and Williams [29], a statistical analysis was performed to examine the relationship between complexity and software security. To assess the relationships between the metrics and vulnerabilities, the researchers employed two correlation coefficients: Pearson correlation coefficient (r) and Spearman rank correlation coefficient ($\rho$). Pearson correlation coefficient assumes a normal distribution of data and measures the strength of linear relationships between variables. On the other hand, the Spearman rank correlation coefficient is a non-parametric test that does not make any assumptions about the distribution of the data and measures the strength of monotonic relationships between variables. Several complexity metrics, including (McCC) McCabe’s cyclomatic complexity [30], Strict cyclomatic complexity, (Nesting) the deepest level of nested control constructs such as “if, while, for, switch, etc.”,(SLOC) the number of source code lines excluding comments, and (Stmt_exe) the number of executable statements, were employed. The results of the analysis revealed the impact of complexity on software security. The study provides empirical evidence that supports the notion that complexity can indeed be considered an enemy of software security.
In a study conducted by [31], the researchers investigated the correlation between software metrics and the presence of security vulnerabilities in C and C++ programs. The analysis revealed that a significant number of the studied metrics exhibited the ability to distinguish between vulnerable and non-vulnerable codes. Some findings aligned with expectations, while others were unexpected. For instance, the vulnerable code was found to contain more comments than the code without vulnerabilities. Additionally, the study highlighted the presence of strong correlations between certain metrics.
Two studies by Chowdhury et al. [32, 33], have identified empirical relationships between software metrics, including cohesion, coupling, complexity, and vulnerability using Spearman rank correlation. These studies categorized the metrics into two groups: design-level metrics and code-level metrics. Design-level metrics can be computed during the design phase, prior to the coding stage. On the other hand, code-level metrics are computed after the coding phase is completed.
  • Code-level complexity metrics include McCabe’s cyclomatic complexity, Nesting, and SLOC. Strict cyclomatic Complexity is another metric that was measured which is identical to cyclomatic complexity except that the AND and OR logical operators are also counted as 1. Comment Ratio is also counted as a complexity measure which is the ratio of the number of comment lines to the number of code lines.
  • Design-level complexity metrics encompass WMC (Weighted Methods per Class), DIT (Depth of Inheritance Tree), and NOC (Number of Children). WMC calculates the number of local methods defined within a class, DIT determines the maximum depth of a class in the inheritance hierarchy, and NOC represents the number of immediate sub-classes or derived classes. CBC is a metric that counts the number of base classes (ancestors)
  • Code-level coupling metrics consist of FanIn and FanOut. FanIn measures the number of inputs used by a function, while FanOut quantifies the number of outputs that are set by a function.
  • Design-level coupling metrics include DIT, NOC, CBC, RFC, and CBO. DIT and NOC are also considered complexity metrics. CBO indicates the level of coupling between object classes. More children in a class mean that it can be coupled to more methods and instance variables. Therefore, NOC and CBC measure coupling as well. RFC indicated the number of methods in the set, including inherited methods.
  • Design-level cohesion metric refers to LOCM (Lack of Cohesion of Methods). LOCM evaluates the cohesion of methods within a class.
The findings of these studies indicate that complexity and coupling metrics in classic software, exhibit a positive correlation with the number of vulnerabilities. This implies that as the complexity and coupling of a software system increase, the likelihood of vulnerabilities also tends to increase. Conversely, cohesion demonstrated a negative correlation with vulnerabilities. This suggests that higher levels of cohesion within a software system are associated with a reduced number of vulnerabilities.
In their research, Shin, Meneely, Williams, and Osborne [34] conducted two empirical studies on two open-source projects to investigate the predictive power of three software metrics (code churn, complexity, developer activity) in identifying vulnerable locations. In their study, the researchers employed logistic regression along with four other classification techniques: decision tree, Random Forest, Naive Bayes, and Bayesian network. The findings revealed that out of the 28 metrics analyzed for both projects, 24 metrics successfully predicted vulnerabilities. Among these 24 metrics, 13 complexity metrics proved to be particularly useful. Examples of these complexity metrics include:
  • CountLineCode: Represents the total number of lines of code in a file.
  • CountDeclFunction: Indicates the number of functions defined in a file.
  • SumCyclomaticStrict: Calculates the sum of the number of conditional statements in a function.
  • SumFanOut: Measures the sum of the number of assignments to parameters when calling a function or modifying global variables.
These complexity metrics played an important role in identifying potential vulnerabilities within the studied projects.
III-E Smart Contract Architecture
Solidity smart contract structures and vulnerabilities exhibit distinct characteristics compared to traditional programming languages. [1]
  • Solidity has a reduced set of data types and libraries compared to most traditional programming languages. It is designed to be simple and deterministic for security and safety reasons.
  • Solidity and other blockchain-specific languages have specific security concerns, such as vulnerabilities related to smart contract bugs, reentrancy, and transaction ordering. Traditional languages also have security concerns, but they differ based on their application domains and execution environments.
  • Solidity programming takes into account the concept of “gas,” which is a measure of the computational resources required to execute a smart contract. Each operation in Solidity consumes gas, and users must pay for the gas when executing a contract. Traditional languages don’t have this gas mechanism as they are not executed on blockchains, and the cost of executing code is typically borne by the server or user’s computing resources.
  • In Solidity, state variables represent the persistent state of a smart contract on the blockchain. They are declared at the contract level and persist across function calls. Traditional languages may have variables with different scopes and lifetimes.
  • Solidity allows the use of function modifiers, which are like reusable code snippets that can be applied to multiple functions in a contract to modify their behavior.
  • In Solidity, functions can be marked as “payable”, indicating that they can receive and handle cryptocurrency payments. Additionally, Solidity has a concept of gas usage for each function, which determines the cost of executing that function on the blockchain.
  • Lastly, Solidity contains new data types such as “address”, “Bytes32”, and “enum”, which further verify that software metrics in Solidity are different.
These distinctive properties of Solidity set it apart from other programming languages, encouraging us to revisit the use of complexity metrics. Although these metrics have been employed in previous studies (mentioned in section 3), the unique nature of smart contracts and the difficulties involved in the Solidity code demand a fresh evaluation of their effectiveness.
In the context of smart contracts, multiple contracts can exist within a single smart contract, resembling the concept of objects in object-oriented programming languages. The diagram depicted in Figure 2 provides an illustration of the structure of a Solidity smart contract.
Figure 2: Structure of a Solidity smart contract
Our objective is to conduct a comprehensive analysis of each contract individually and identify vulnerable contracts within the smart contract ecosystem. By focusing on each contract separately, we can gain insights into their specific vulnerabilities and assess their overall security. This approach allows us to pinpoint and address vulnerabilities at the contract level, enabling more targeted and effective security measures.
IV Complexity Metrics
The complexity metrics we employ in our analysis are obtained through the use of Solmet [35], a static analysis tool specifically designed for extracting complexity metrics from Solidity code. Solmet enables us to extract a comprehensive set of 21 complexity metrics for each contract individually. Table II provides a detailed overview of all the metrics extracted by Solmet.
  • SLOC: The Source Lines of Code (SLOC) metric captures the total number of lines in a software artifact, including comments and white spaces. This metric was measured in [34] as CountLineCode.
  • LLOC: Logical Lines of Code (LLOC) is a complexity metric that measures the number of executable lines of source code in a software artifact (lines without comments and white spaces). While LLOC primarily indicates the size of the codebase, prior research has shown a strong correlation between SLOC and complexity. This metric was named SLOC in other papers [29, 32, 33]
  • CLOC: The number of comment lines (CLOC) metric, as utilized in prior studies [32, 33], plays a role in code-level complexity metrics by representing the comment ratio in software code. The comment ratio is calculated by dividing the number of comment lines by the number of code lines. By considering both CLOC and the number of lines of code (LLOC), the comment ratio can be determined, providing insights into the proportion of comments relative to the actual code. According to the findings in [36], it is suggested that highly complex program units tend to have a higher comment ratio, indicating a larger number of comments associated with each line of code. This suggests that the comment ratio may provide some insights into the complexity of the code.
  • NF: The number of functions (NF) metric was evaluated and demonstrated its usefulness in the study conducted by Shin et al.  [34]. This metric focuses on quantifying the total number of functions within a contract. This was called RFC in [32, 33]. The weighted version of this metric was also used in their work as WMC.
  • WMC The weighted sum of McCabe’s style complexity (WMC) metric, which considers the complexity of individual functions within a contract, has been discussed in previous studies such as Shin et al. [29, 34], Chowdhury et al. [32, 33]. WMC is an extension of the strict cyclomatic complexity metric that incorporates weights assigned to each function.
  • NL: NL was also measured in [29, 32, 33] as Nesting, is the deepest nesting level of control structures in functions summed for a contract. The presence of highly nested control structures (such as if, while, for, switch, etc.) in a function can contribute to the complexity of a program entity. When control structures are deeply nested, it can make the code more intricate and challenging for programmers to comprehend.
  • NLE: The Nesting Level of Else-If (NLE) metric specifically focuses on the nesting level of “else-if” control structures within the code. It is similar to the Nesting Level (NL) metric, but it is only counted for “if” and “else-if” control structures.
  • NUMPAR: The NUMPAR metric, also known as the Number of Parameters, refers to the count of parameters or inputs that functions within a contract have.
  • NOS: The number of statements (NOS) was also referenced as stmt_exe in the study conducted by Shin et al. [29]. NOS counts conditional and executable statements in a contract.
  • DIT: The depth of inheritance tree (DIT) was also discussed as a design-level complexity and coupling metric in the paper referenced [32, 33]. The maximum depth of a class in the inheritance tree, also known as the Depth of Inheritance Tree (DIT), refers to how deep a class is positioned within the hierarchy of inherited classes. A higher DIT value indicates that the class inherits a greater number of methods, which can contribute to increased complexity in predicting its behavior. The deeper a class is in the inheritance hierarchy, the more complex it becomes to understand and anticipate its functionality [37].
  • NOA & NOD: In the research papers by Chowdhury et al.[32, 33], the metrics “number of ancestors” (NOA) and “number of descendants” (NOD) were used and referred to as CBC and NOC, respectively. These metrics are related to the inheritance relationships among classes in an object-oriented system.
  • CBO: The Coupling Between Object classes (CBO) is a metric that measures the number of other classes coupled to a specific class (or contract). It indicates the level of dependency and interaction between classes within an object-oriented system. A higher value of CBO suggests that a class is more tightly coupled to other classes, indicating a potentially higher complexity and interdependencies. Chowdhury et al. [32, 33] conducted research where they measured the CBO metric and established its relationship with vulnerabilities.
  • NA: The metric NA, which stands for Number of Attributes, represents the count of attributes associated with a class or a contract in Solidity. Attributes, in the context of Solidity contracts, refer to various modifiers or keywords that are added before a variable name to define its characteristics or structure. Examples of attributes in Solidity contracts include “external,” “indexed”, “internal,” “payable” and others. This metric is counting the vocabulary size of a contract.
  • NOI: The metric NOI (Number of Outgoing Invocations), also referred to as fan-out, quantifies the count of outgoing invocations or function calls from a specific function or method. It represents the number of other functions or methods that are called within the target function. The fan-out metric measures the degree of coupling or dependencies between the target function and other functions in the codebase. It indicates how extensively the target function relies on and interacts with other functions. A higher fan-out value suggests that the target function invokes or depends on a larger number of external functions. Fan-out was also previously measured in [34, 32, 33].
  • Avg. McCC: The metric Avg. McCC (Average McCabe’s Cyclomatic Complexity) is a measure of the average cyclomatic complexity of a software artifact, such as a function, module, or class. It quantifies the complexity of the control flow within the artifact based on the number of decision points and possible paths through the code. McCabe’s cyclomatic complexity is calculated by counting the number of independent paths through the code, which is determined by the number of decision points, such as conditionals (if statements, switch cases), loops (for, while, do-while), and logical operators. The higher the cyclomatic complexity, the more complex and intricate the control flow of the code. The McCC metric has been mentioned in the previously cited papers as a relevant measure of complexity and vulnerability.
  • Avg. : Each of the metrics, namely Avg. NL (Average Nesting Level), Avg. NLE (Average Nesting Level of Else-If), Avg. NUMPAR (Average Number of Parameters), Avg. NOS (Average Number of Statements), and Avg. NOI (Average Number of Outgoing Invocations), calculates the sum of its respective metric for each contract. This sum is then divided by the number of functions present in the contract to obtain the average value.
Each metric that counts a number, such as the number of lines of code or the number of parameters, represents either the vocabulary size or the code size of a contract. These metrics have been highlighted in a prior study conducted by Peitek et al. [38] as factors that require more attention and mental effort from programmers. As a result, they can increase cognitive load during code comprehension and impose a greater burden on the working memory of programmers. Consequently, these metrics can serve as indicators of potential vulnerabilities in a contract. By examining these metrics, we can gain insights into the complexity and potential risks associated with the code.
TABLE II: Solidity complexity features analyzed by Solmet
SLOC number of source code lines
LLOC number of logical code lines (lines without empty and comment lines)
CLOC number of comment lines
NF number of functions
WMC weighted sum of McCabe’s style complexity over the functions of a contract
NL the deepest nesting level of control structures in functions summed for a contract
NLE nesting level else-if
NUMPAR number of parameters
NOS number of statements
DIT depth of inheritance tree
NOA number of ancestors
NOD number of descendants
CBO coupling between object classes
NA number of attributes (i.e. states)
NOI number of outgoing invocations (i.e. fan-out)
Avg. McCC McCabe’s cyclomatic complexity
Avg. NL Average number of NL
Avg. NLE Average number of nesting level else-if
Avg. NUMPAR Average number of parameters
Avg. NOS Average number of statements
Avg. NOI Average number of outgoing invocations
V Dataset and Labeling
For our research, we utilized the dataset provided by Liu et al [39]. They made their work publicly available on their GitHub repository, enabling further investigation. We use the “Resource 3” on their GitHub [^2] https://github.com/Messi-Q/Smart-Contract-Dataset, which contains 8 types of vulnerabilities. The vulnerability types they measured can be found in Table III. This newly published dataset, released in 2023, includes 2,953 smart contracts and a total of 16,239 contracts which contained 258 vulnerable and 15981 neutral contracts.
TABLE III: Vulnerabilities labeled in the dataset
Vulnerability Type
Integer Underflow/Overflow
Dangerous Ether Strict Equality
Reentrancy
Timestamp Dependency
Block Number Dependency
Dangerous Delegatecall
Ether Frozen
Unchecked External Call
VI Experimental Setup
To answer our questions, we employ several statistical techniques and machine-learning supervised classifiers.
VI-A Statistical Techniques
Spearman’s Correlation Coefficient: The Spearman correlation coefficient is a statistical measure that allows us to assess the monotonic relevance of each metric separately to the presence of vulnerabilities in smart contracts [40]. It determines the strength and direction of the monotonic relationship between two variables and is a non-parametric method, which means it does not assume a specific distribution for the variables. This statistical measure is particularly useful when analyzing data without relying on the magnitudes of the values. Instead, it focuses on the ranks or order of the values. This property makes it robust and less sensitive to outliers in the data [41]. By considering the ranks, the Spearman rank correlation can capture non-linear relationships between variables, making it a suitable choice for assessing the association between complexity metrics and vulnerability indicators [32]. The formula for calculating the Spearman correlation coefficient ($\rho$) is:
$\rho=\frac{1-(6*sum)}{(n*(n^{2}-1))}$ (1)
In this formula, ’n’ represents the number of observations, and ’sum’ represents the sum of squared differences. The formula calculates the sum of squared differences between the ranks of corresponding observations for the two variables. It then divides this sum by a scaling factor to obtain the Spearman correlation coefficient ($\rho$). The coefficient ranges between -1 and 1, where values close to 1 indicate a strong positive monotonic relationship, values close to -1 indicate a strong negative monotonic relationship and values close to 0 indicate a weak or no monotonic relationship.
In statistical analysis, the significance of a correlation is commonly evaluated using a p-value, which represents the probability of observing a correlation by chance. A smaller p-value indicates a higher level of confidence in the significance of the correlation. The threshold commonly used to determine statistical significance is 0.05, which corresponds to a 95% confidence level. If the p-value is less than or equal to 0.05, it is considered statistically significant, indicating that the observed correlation is unlikely to occur by chance [41]
Paired T-test: A paired t-test, also known as a dependent t-test or a matched-pairs t-test, is a statistical hypothesis test used to compare the means of two related groups [42]. It is designed for situations where you have two sets of measurements on the same subjects or items, and the values in each set are paired together. The pairing can occur because the measurements are taken from the same individual or item under different conditions or at different time points.
The paired t-test is used to determine if there is a significant difference between the means of the two sets of paired data. The assumption behind this test is that the differences between the paired data points are normally distributed.
In this test, If the p-value is smaller than the chosen significance level (in our case 0.05), you reject the null hypothesis and conclude that there is a significant difference between the means of the two paired groups [42].
Confidence Interval: The primary objective of computing a statistic from a random sample is to approximate the mean of the entire population. However, the accuracy of this approximation is always a concern. To address this issue, a confidence interval is employed, which provides a range of values likely to contain the true population parameter of interest. The confidence level for the interval is chosen by the specialist, typically set at 95%. This means that if the population is sampled multiple times, the resulting interval estimates would encompass the true population parameter in approximately 95% of cases [42]. The confidence interval provides a valuable tool to gauge the reliability and precision of the sample statistic in estimating the characteristics of the entire population.
VI-B Machine Learning Classifiers
To assess the predictive capability of Machine Learning models using these complexity metrics for vulnerable contract identification and answer question 5, we employed various ML classifiers along with Multilayer Perceptron. The evaluation is conducted using 10-fold cross-validation on our dataset, ensuring the robustness and reliability of the results. We report four commonly used evaluation metrics: accuracy, precision, recall, and F1-measure. In the evaluation of our classifiers, we utilize the basic criteria of True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN). These base values serve as the foundation for calculating other performance indicators.
  • A True Positive (TP) occurs when the classifier correctly identifies a positive instance (vulnerable contract) as positive. It represents a correct prediction of vulnerability.
  • A True Negative (TN) is the correct identification of a negative instance (non-vulnerable contract) as negative. It signifies a correct prediction of non-vulnerability.
  • On the other hand, a False Positive (FP) happens when the classifier incorrectly classifies a negative instance as positive (non-vulnerable contract as vulnerable).
  • Lastly, a False Negative (FN) occurs when the classifier misclassifies a positive instance as negative (vulnerable contract as non-vulnerable).
The relationship between these four criteria is summarized in Table IV.
TABLE IV: Basic evaluation criteria
Criteria
TP
Correct Result
FP
Unexpected result
FN
Missing result
TN
Correct absence of result
In our experimental results, we report several performance metrics to evaluate the effectiveness of our models. Similar to MODNN, we calculate the “weighted average” scores of these metrics to account for the class imbalance. The “weighted average” means that the scores are calculated by considering the class distribution and giving more weight to the underrepresented class (vulnerable contracts) in the evaluation.
The performance metrics we report are as follows:
  • Accuracy: It measures the fraction of correct predictions made by our models. It is calculated as the sum of true positives and true negatives divided by the total number of predictions.
    $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (2)
  • Recall: Also known as sensitivity or true positive rate, indicates how many of the vulnerable contracts our models could find. It is calculated as the ratio of true positives to the sum of true positives and false negatives.
    $Recall=\frac{TP}{TP+FN}$ (3)
  • Precision: Reports how many of the vulnerable contracts that our models could find were indeed vulnerable. It is calculated as the ratio of true positives to the sum of true positives and false positives.
    $Precision=\frac{TP}{TP+FP}$ (4)
  • F1 score: It is the harmonic mean of precision and recall, providing a balanced measure of the model’s performance. It is calculated using the formula:
    $F1=\frac{2*Precision*Recall}{Precision+Recall}$ (5)
  • FPR: The False Positive Rate and False Negative Rate are two important metrics used to evaluate the performance of binary classification models, such as ours. These rates help us understand how well a model is able to distinguish between the two classes. The False Positive Rate is the ratio of the number of false positives to the total number of actual negatives (true negatives + false positives). In this context, FPR refers to the proportion of neutral codes that are incorrectly classified as vulnerable. It is calculated by:
    $FPR=\frac{FP}{TN+FP}$ (6)
  • FNR: The False Negative Rate is the ratio of the number of false negatives to the total number of actual positives (true positives + false negatives). In this context, FNR refers to the proportion of vulnerable codes that are incorrectly classified as neutral. FNR is calculated by:
    $FPR=\frac{FN}{TP+FN}$ (7)
Some popular Machine learning classification models are used in this context. For classification, we utilize the following models:
  • K-Nearest Neighbors (KNN) with K values of 1, 3, and 5. Higher values of K are not considered to prevent underfitting.
  • Decision Tree classifier based on entropy.
  • AdaBoost classifier with 300 estimators.
  • Random Forest classifier with 100 estimators.
  • Multi-Layer Perceptron classifier
All models are implemented using Scikit-learn [43], a popular machine-learning library in Python.
VI-C Managing the Dataset
To address the data imbalance, we apply class weighting during the model-tuning process and weighted averaging while reporting the results. This ensures that the models can effectively handle the imbalanced nature of our dataset and produce more accurate results.
Alternative approaches for handling imbalanced datasets, such as SMOTE (Synthetic Minority Over-sampling Technique) [44], might still be necessary. SMOTE addresses class imbalance by combining oversampling and undersampling methods. Oversampling involves generating synthetic samples for the minority class by interpolating between existing instances, effectively increasing the number of minority class samples. However, this method is not suitable for our work, as it would generate an artificial dataset with metric values that may never be seen in real code. Conversely, undersampling involves randomly removing instances from the majority class to reduce its dominance. This method may be useful since it does not create artificial metric values. By balancing the dataset through these combined techniques, SMOTE can improve model performance on imbalanced datasets.
To provide a comprehensive evaluation, we report the average classification results over the ten folds, along with the corresponding standard deviation. This allows us to assess the consistency and stability of the models’ performance across different subsets of the dataset.
For our execution environment, we utilized Google Colab  [^3] https://colab.research.google.com/ in its free-plan hardware accessibility mode, on its CPU. Google Colab is a cloud-based platform used to execute our programs and conduct experiments. Google Colab provides a convenient environment for running Python code, leveraging its powerful computational resources and integration with popular libraries and frameworks such as Scikit-learn.
VII Experimental Results
In this results section, we give answers to each research question asked.
RQ1: Are there correlations among complexity metrics? Our objective was to identify potential redundancies or inconsistencies among the various complexity metrics available. To address this question, we employed Spearman’s correlation coefficient. Initially, we divided the codes into two groups: vulnerable codes and neutral codes (without vulnerabilities). Subsequently, we calculated Spearman’s correlation coefficient for each metric within each group. This approach enabled us to identify any potential redundancies among the metrics in both code groups. For our analysis, we considered two variables to be strongly correlated if the coefficient ($\rho$) exceeded 0.9 [40]. Figure 3 provides a visual representation of the correlation analysis, distinguishing between vulnerable (blue) and neutral (green) codes. The results of this analysis give information about the interrelationships among the complexity metrics within each code group, helping us understand which metrics are closely related to each other. The highlighted squares in the correlation matrix demonstrate an unusually strong association between certain metrics, suggesting potential redundancy among them.
Figure 3: Correlation between metrics in vulnerable contracts (blue) and neutral contracts (green)
Among all the coefficients calculated for each group, the p-values for the neutral group were consistently below 0.05, indicating a high level of confidence in the results. However, exceptions were noted in the relationships between CLOC and NOD, DIT and NOD, and NOA and NOD, with p-values of 0.50, 0.09, and 0.09, respectively. Unfortunately, for the vulnerable group, the p-values were higher than 0.05 in many instances. Given the high number of cases, we only report those that are important for our use, specifically when a high correlation is observed. For a full report, you can look at our implementation’s source code which is publicly available at https://zenodo.org/records/13364187
The potential redundant metrics identified in the analysis are as follows:
  • SLOC: Strongly related to NOI, NOS, WMC, NF, and LLOC. SLOC counts the number of source lines of code, including comments and empty spaces.
  • NF and WMC: Conceptually unrelated, but show a relationship. NF counts the number of functions, and WMC calculates McCabe’s cyclomatic complexity over functions in a contract. The similarity could be due to the relatively lower number of control structures (to lower gas prices) in smart contracts, resulting in equal numbers for functions and McCabe’s complexity. Furthermore, both show a high correlation with similar metrics such as NOI, NOS, NUMPAR, NLE, and NL.
  • NLE, NL, Avg. NLE, and Avg. NL: These metrics measure the nesting level of control structures. They are equal due to the limited use of “if-else” and “loop” structures in smart contracts, suggesting that only one of these metrics may be useful.
  • Avg. McCC, Avg. NLE, and Avg. NL: Similarity due to the same reasons as mentioned before, indicating potential redundancy.
  • NOA and DIT: These metrics, derived from the inheritance tree, exhibit similarity, possibly indicating redundant information.
  • NOI and Avg. NOI: Are equal due to the limited use of invocations in smart contracts, suggesting that only one metric might be necessary. The high correlation between NOI and other metrics of SLOC, NF, WMC, and NOS was mentioned previously.
Fortunately, all the p-values between these metrics with high correlation were less than 0.05, thus this report is reliable. Considering these potential redundancies, it is crucial to carefully select and include only the most relevant and distinct complexity metrics in the vulnerability prediction models. Removing redundant metrics can optimize the feature set and improve the efficiency and interpretability of the models while achieving comparable predictive performance.
RQ2: Is there a correlation between each complexity metric and the existence of vulnerability? We aimed to determine whether certain complexity metrics are associated with the presence of vulnerabilities in individual contracts. This part also utilizes the Spearman Correlation Coefficient but without separately grouping the codes into vulnerable and neutral. In our study, all of the calculated p-values for the correlations between complexity metrics and vulnerability indicators were less than 0.05. The result of the Spearman correlation coefficient for each metric is represented in Table V. To analyze the result more comprehensively, figure 4 illustrates the result in a bar plot.
Figure 4: Bar plot of Spearman correlations between complexity metrics and vulnerabilities
TABLE V: Correlations between complexity metrics and existence of Vulnerabilities
Metrics Correlation Coefficient P-Value
SLOC 0.153 <0.05
LLOC 0.132 <0.05
CLOC -0.077 <0.05
NF 0.123 <0.05
WMC 0.130 <0.05
NL 0.146 <0.05
NLE 0.144 <0.05
NUMPAR 0.082 <0.05
NOS 0.160 <0.05
DIT 0.099 <0.05
NOA 0.102 <0.05
NOD -0.042 <0.05
CBO 0.182 <0.05
NA 0.170 <0.05
NOI 0.165 <0.05
Avg. McCC 0.127 <0.05
Avg. NL 0.121 <0.05
Avg. NLE 0.116 <0.05
Avg. NUMPAR -0.037 <0.05
Avg. NOS 0.144 <0.05
Avg. NOI 0.145 <0.05
The interpretation of the correlation coefficient depends on the context of its usage. According to Cohen et al., [41], the following guidelines can be used to interpret the strength of the relationship:
  • A correlation coefficient of less than 0.3 indicates a weak correlation.
  • A correlation coefficient between 0.3 and 0.5 suggests a medium correlation.
  • A correlation coefficient greater than 0.5 indicates a strong correlation.
These guidelines provide a qualitative assessment of the strength of the relationship between the variables being studied. In our analysis, we apply these interpretations to understand the degree of association between complexity metrics and vulnerability in Solidity smart contracts. Our analysis reveals that among the complexity metrics used in this study, all metrics show weak correlations with vulnerability in Solidity smart contracts. This finding suggests that no single metric alone can effectively distinguish between vulnerable and neutral codes in smart contracts, which contrasts with previous research on traditional programming languages [31, 32] showing stronger correlations between certain complexity metrics and vulnerability, often with coefficients exceeding 0.5.
Table VI illustrates the comparison of correlation outcomes between complexity metrics in smart contracts and the same metrics in traditional programming languages. Only the metrics consistent with previous research are included. The table findings indicate that while in other languages these metrics exhibited a monotonic relation with vulnerability presence, this relationship was not consistently observed in the context of smart contracts.
TABLE VI: Comparing the correlation values of complexity and existence of vulnerability between smart contracts (SC) and traditional programs (TP)
Metrics SC Corr TP Corr
SLOC 0.153 0.292
LLOC 0.132 0.288
CLOC -0.077 0.324
NF 0.123 0.434
WMC 0.130 0.429
NL 0.146 0.532
NLE 0.144 -
NUMPAR 0.082 -
NOS 0.160 0.292
DIT 0.099 0.459
NOA 0.102 -
NOD -0.042 0.642
CBO 0.182 0.454
NA 0.170 -
NOI 0.165 0.514
Avg. McCC 0.127 -
Avg. NL 0.121 -
Avg. NLE 0.116 -
Avg. NUMPAR -0.037 -
Avg. NOS 0.144 -
Avg. NOI 0.145 -
RQ3: Are the metrics different in vulnerable and neutral codes? We sought to investigate whether the complexity metrics can effectively distinguish between functions with reported vulnerabilities and those without reported vulnerabilities. To assess whether the metrics in each group (vulnerable and neutral) can effectively discriminate between vulnerable and neutral codes, we employed the paired t-test. This statistical test was conducted by comparing the metric values of the vulnerable group with their corresponding values in the neutral group. A p-value greater than 0.05 indicates that the null hypothesis is true, suggesting that the metric’s values are derived from the same distribution and cannot effectively discriminate between vulnerable and neutral codes. On the other hand, a p-value less than or equal to 0.05 suggests that there are significant differences in the metric values between the two groups, indicating the potential of the metric to discriminate vulnerable codes from neutral codes. The p-values and the results are available in Table VII
TABLE VII: The paired t-test between each metric’s values in vulnerable and its corresponding values in neutral codes
Metrics Discriminative P-Value
SLOC yes <0.05
LLOC yes <0.05
CLOC yes <0.05
NF yes <0.05
WMC yes <0.05
NL yes <0.05
NLE yes <0.05
NUMPAR yes <0.05
NOS yes <0.05
DIT yes <0.05
NOA yes <0.05
NOD yes <0.05
CBO yes <0.05
NA yes <0.05
NOI yes <0.05
Avg. McCC yes <0.05
Avg. NL yes <0.05
Avg. NLE yes <0.05
Avg. NUMPAR yes <0.05
Avg. NOS yes <0.05
Avg. NOI yes <0.05
The table shows that all metrics can be sufficiently discriminative in distinguishing vulnerable codes from neutral codes in Solidity smart contracts.
RQ4: How different are the metrics in vulnerable codes compared to neutral ones? We intended to examine whether higher complexity in code is linked to a higher likelihood of vulnerability occurrence. This approach enables a more targeted and efficient allocation of resources toward securing smart contracts.
To answer this question, we need to compare the mean of each group. To compare the means of each group, we employed confidence intervals, ensuring that the resulting intervals encompass the true population parameter in around 95% of cases. This statistical approach allowed us to effectively evaluate the differences between vulnerable and neutral codes in Solidity smart contracts. The comparison results are presented in Figure 5.
Figure 5: Confidence interval comparison in vulnerable contracts (1) and neutral contracts (0)
In the figure, it can be observed that, apart from the three exceptions, every complexity metric is higher in vulnerable contracts, indicating a positive correlation between complexity and vulnerability. However, the three metrics Avg. NUMPAR, CLOC, and NOD contrast this trend. These findings are distinct from those reported in [31], demonstrating the unique characteristics and complexities present in Solidity smart contracts, especially in the case of CLOC, which measures the number of lines of comments in the code. In previous work, this metric showed higher values in vulnerable codes. However, in our study, higher CLOC values were observed in neutral codes. This suggests that writing comments in smart contracts might help identify and mitigate vulnerabilities within the code.
VIII Discussion
The dataset used is newly created but only comprises 16,239 contracts, which might be considered small for our 21 metrics. Although there are several large smart contract datasets containing new source codes [45], a sufficiently large and new dataset labeled with vulnerabilities has yet to be developed. Given these limitations, caution must be exercised when generalizing the results of this dataset to the current state of Solidity and its vulnerabilities. Future research endeavors should strive to include larger datasets, up-to-date analysis tools, and labeling approaches that align with the latest versions of Solidity. By doing so, we can ensure that our conclusions and evaluations accurately reflect the most relevant and current vulnerabilities present in Smart Contracts.
In future work, we have identified several areas for improvement and expansion in our study. One of our objectives is to include additional complexity metrics that are specific to smart contracts, such as the amount of gas expended, to enhance our analysis.
Furthermore, we plan to explore the inclusion of comprehension complexity metrics, such as cognitive complexity [46, 47], in our evaluation. Cognitive complexity measures the difficulty of understanding code by considering factors such as nested conditions, control flow, and human cognition. By including such metrics, we can gain a better understanding of the cognitive load imposed on developers and its application in vulnerabilities of Smart Contracts. These future efforts will contribute to a more comprehensive and robust assessment of Smart Contract security.
IX Conclusion
Our study examined the use of complexity metrics as indicators of vulnerable code locations in Solidity smart contracts. We found that vulnerable functions exhibit distinct characteristics in terms of code complexity compared to non-vulnerable functions.
Throughout our investigation, we have obtained a valuable understanding of the relationship between complexity metrics and vulnerability detection in Solidity smart contracts. From our findings, we can draw the following conclusions:
  1. 1.
    Some complexity metrics are highly correlated and redundant, suggesting that certain metrics may not provide additional information beyond what others already capture.
  2. 2.
    Individual complexity metrics alone are insufficient to predict vulnerable contracts effectively, as their correlations with the existence of vulnerabilities are weak.
  3. 3.
    Complexity metrics exhibit distinct values and distributions in vulnerable and non-vulnerable codes, demonstrating their discriminative power.
  4. 4.
    Most complexity metrics tend to have higher values in vulnerable codes, except for a few specific ones, such as the number of comment lines, number of descendants, and average number of parameters, which were lower in vulnerable codes.
Although the Spearman rank correlation coefficients indicated low individual correlations between each complexity metric and vulnerability, our machine learning models still achieved good results. This suggests that no single metric can effectively distinguish between vulnerable and invulnerable contracts on its own. However, when combined, the metrics collectively provide practical information and enhance the performance of the models. This highlights the importance of considering multiple complexity metrics together rather than relying on a single metric.
This study provides practical evidence of the effectiveness and relevance of complexity metrics in predicting vulnerabilities in Solidity smart contracts and analyses the power of each metric to help researchers understand their individual influence. These metrics can enhance existing models and methodologies by providing additional insights into code vulnerability assessment.
Data Availability
The implementations, source code, data, and experimental results are publicly available at https://zenodo.org/records/13364187